OpenStudy (anonymous):

x^^2 + 10x + 24 = 0 by completing the square

OpenStudy (anonymous):

x^2+10x +24=0 consider x^2+10x first x^2 +10x = x^2 +(1/2)(20)x = x^2 +(2)(5x)+5^2-5^2 = (x+5)^2-5^2 x^2+10x +24=0 (x+5)^2-5^2+24=0 (x+5)^2=1 (x+5) = 1 or -1 x=-4 or -6